OpenStudy (sammyalabamy):

What is the molarity of a 0.25-liter solution containing 52 grams of HBr?

OpenStudy (photon336):

\[Molarity = \frac{ Moles }{ Liter }\]

OpenStudy (sammyalabamy):

thank you @Photon336

OpenStudy (photon336):

okay so first step we need to find the molar mass HBr

OpenStudy (sammyalabamy):

80.912

OpenStudy (photon336):

okay so now we know the molar mass of Hbr and we know how many grams we have. we need to find the number of moles like this.. \[52~grams~HBr*(\frac{ mole }{ 81grams }) = moles~HBr\]

OpenStudy (sammyalabamy):

0.6419

OpenStudy (photon336):

okay now we know 0.6419 moles of HCL and our volume is in liters 0.25L so now we have all we need to complete this. \[\frac{ 0.6419~moles }{ 0.25~L } = Molarity \]

OpenStudy (sammyalabamy):

2.5676 Thank you so much @Photon336

OpenStudy (photon336):

no problem you could round up to 2.6 mol/L or 2.6 M M is the symbol for molarity btw

OpenStudy (photon336):

Just make sure that if you're given volume it is in liters.